# The Vital Role of Gradient Clipping in Byzantine-Resilient Distributed Learning

### v3

@@ 4 Empirical Evaluation
We show in Table 1, and Figures 2a and 4, the metric of worst-case maximal accuracy. For each of the five Byzantine attacks executed (see Appendix D.3), we record the maximal accuracy achieved by Robust-DSGD during the learning procedure under that attack. The worst-case maximal accuracy is thus the smallest maximal accuracy reached across the five attacks. As the attack cannot be known in advance, this metric is critical to accurately evaluate the robustness of aggregation methods, since it provides an estimate of the potential worst-case performance of the algorithm. Furthermore, we use the Dirichlet (Hsu et al., 2019a) distribution of parameter $\alpha$ to simulate data heterogeneity. The comprehensive experimental setup can be found in Appendix D. In this section, we focus on presenting our results for the CWTM and GM aggregation methods applied to MNIST and CIFAR-10. Results for Fashion-MNIST, as well as for the CWMed and MK aggregations, are provided in Appendix E. All aggregation methods in our experiments are pre-composed with NNM (Allouah et al., 2023a), but we omit explicit mention of NNM in their names throughout the text for simplicity.
@@
ARC boosts robustness in high heterogeneity. Figure 3 shows the performance of Robust-DSGD against the FOE attack when using ARC opposed to no clipping. In low heterogeneity (i.e., $\alpha=0.5$, see Figure 7), the performances of ARC and no clipping are comparable for both aggregations. When the heterogeneity increases ($\alpha=0.1$), the benefits of using ARC are more visible. CWTM and GM significantly struggle to learn (with a very large variance), while the same aggregations composed with ARC almost match the accuracy of DSGD towards the end of the training. In extreme heterogeneity, the improvement induced by our method is the most pronounced, as ARC enables both aggregations to reach a final accuracy close to 90%. Contrastingly, the same aggregations without clipping stagnate at around 10% throughout the training. Interestingly, only $f=1$ adversarial worker among $n=11$ (i.e., less than 10%) suffices to completely deteriorate the learning when no clipping is applied, highlighting the necessity of ARC in extreme heterogeneity.
Varying heterogeneity for fixed $f$. Figure 4a compares the worst-case maximal accuracies achieved by Robust-DSGD when $f=1$, with ARC and without, for varying levels of heterogeneity. While CWTM and GM yield comparable accuracies with ARC in low heterogeneity ($\alpha\geq 0.5$), their performance significantly degrades when $\alpha$ drops below that threshold. Indeed, when $\alpha=0.1$, their accuracies drop below 65%, while ARC enables the same aggregations to maintain their accuracy at just below 98%. In extreme heterogeneity, the performance of the aggregations without clipping completely deteriorates with accuracies close to 15%. Contrastingly, ARC efficiently mitigates the Byzantine attacks, resulting in accuracies above 85% in the worst case for both aggregations. Similar plots for $f=3,5,7$ convey similar observations (see Appendix E.1). Essentially, in low heterogeneity or when the fraction of adversarial workers is small, no clipping yields good empirical results and ARC performs at least as well as no clipping. However, ARC induces a significant improvement near the breakdown point of SOTA robust aggregation schemes (e.g., high heterogeneity or large $f$).
ARC increases the breakdown point in adverse settings. Figure 2a of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%. This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity. However, composing them with ARC increases the breakdown point of these aggregations to $f=3$. Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case. However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%. Moreover, even when the heterogeneity is not large, ARC still produces a significant improvement when the fraction of adversarial workers increases in the system. Indeed, in Figure 4b , the performances of ARC and no clipping are comparable for $f\leq 3$. However, the improvement induced by ARC is much more visible when $f\geq 5$. Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping . This observation indicates the raise of the breakdown point due to ARC in practice. Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix E.1.
Figure 5: FOE on heterogeneous CIFAR-10 with $\alpha=0.075$. $f=1$ among $n=17$.
Improved robustness on CIFAR-10. We also conduct experiments on the more challenging CIFAR-10 task with $n=17$ and $f=1$. Table 2 shows the worst-case maximal accuracies achieved by ARC and no clipping for four robust aggregations. For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for all aggregations, while the same aggregations without clipping yield lower accuracies (with a larger variance across seeds). For instance, GM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC. In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance). However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM and CWMed achieving 16% and 13.7% in worst-case accuracy, respectively. This can also be seen in Figure 5, where FOE completely degrades the learning when ARC is not used.
@@ 5.1 Improvement of convergence guarantees
In this section, we characterize the improvement of Robust-DGD with ARC over the lower bound recalled in Lemma 2.3. Specifically, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$. First, we establish in Lemma 5.1 a key property of ARC, which is crucial for deriving the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
Lemma 5.1 (Bounded Aggregation Output).
Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method. For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, the following holds true:
$\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$ (1)
@@
In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in 2.3. For ease of presentation, we denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, let
$\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$ $\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right),$
@@
Influence of model initialization and data heterogeneity. Note that the smaller the bound on the initial gradients (i.e., the better the model initialization), the greater the improvement induced by ARC. Specifically, a decrease in $\zeta$ leads to a reduction in $\rho$, which subsequently lowers $\Psi(G,B,\rho)$. As a result, for a fixed value of $\xi$ (i.e., for a fixed ratio of adversarial workers), the condition $\xi\leq\boldsymbol{\upsilon}/\Psi(G,B,\rho)$ is satisfied for a smaller $\boldsymbol{\upsilon}$, thereby yielding a lower stationarity error. This theoretical deduction is empirically validated in Section 5.2. A similar improvement occurs when increasing data heterogeneity, while keeping all other factors unchanged. Specifically, as $G$ increases, $\Psi(G,B,\rho)$ decreases, allowing for a smaller reduction factor $\boldsymbol{\upsilon}$ and thus a larger improvement in performance. This trend is also validated empirically in Figures 3 and 4a.
Influence of $\boldsymbol{f/n}$. Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 2a and 4b. Indeed, we show in Theorem C.4 (a more complete version of Theorem 5.2) that ARC effectively increases the $\mathsf{BP}$ of ($f,\kappa$)-robust methods from $\frac{1}{2+B^{2}}$ to $\frac{1}{2}$, provided the initial honest gradients are bounded in norm. Lastly, we would like to recall that when the workers’ gradients at model initialization are arbitrarily large, the convergence guarantee of Robust-DGD with ARC reduces to that of classic Robust-DGD (see Section 3). In other words, there is no downside to incorporating ARC.
@@ 5.2 Influence of model initialization on empirical robustness
Figure 6: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, both with and without ARC, on MNIST with $\alpha=0.1$, 10 honest worker, and $f=1$ Byzantine worker. The x-axis represents settings with progressively worse model initialization. This figure complements Figure 2b.
Figures 2b and 6 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers. We investigate two regimes of heterogeneity: extreme heterogeneity and $\alpha=0.1$, and consider $f=1$ adversarial worker. The experiment proceeds as follows: we begin with the default model parameters initialized by PyTorch (i.e., the same ones used in Section 4), which represent a well-chosen set of initial parameters. These parameters are then scaled multiplicatively by a factor $\mu$ where larger values of $\mu$ correspond to progressively worse initialization, and vary $\mu\in\{1,...,5\}$. The results show that, under well-initialized conditions ($\mu=1$), ARC significantly enhances the performance of Robust-DSGD, achieving a substantial improvement in worst-case maximal accuracy, particularly under extreme heterogeneity. In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 2b). This ($\mu=1$) corresponds to the initialization conditions of the empirical results presented in Section 4. As $\mu$ increases (i.e., the initialization worsens), the performance of ARC-enhanced Robust-DSGD gradually declines, with noticeable degradation starting from $\mu=4$. Nevertheless, even at this point, ARC still offers a performance advantage over Robust-DSGD without clipping. By $\mu=5$, the performance of Robust-DSGD with ARC converges to that of plain Robust-DSGD, both achieving an accuracy of around 10%. Another key observation from Figures 2b and 6 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD. Both exhibit similarly poor performance when $\mu=5$, and their accuracies are comparable when $\mu\leq 3$, especially when $\alpha=0.1$. This suggests that ARC is particularly effective at exploiting good model initialization, similar to the performance of DSGD in the absence of Byzantine workers. In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 2b. These findings highlight the important influence of model initialization on the robustness of aggregation methods, and empirically validate our theoretical findings in Section 5.1.
@@ 6 Related Work
Gradient clipping is a well-known technique for tackling exploding gradients in deep learning (Goodfellow et al., 2016). It has been extensively analyzed in the centralized setting (Zhang et al., 2020b; a; Koloskova et al., 2023), with applications to differential privacy (Pichapati et al., 2019). Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise Gorbunov et al. (2020; 2022); Danilova (2023). The considered proof techniques, however, do not apply directly to our adversarial distributed learning setting. A similar study on the impact of clipping also exists for distributed settings (Zhang et al., 2022; Khirirat et al., 2023).
@@
While adaptive clipping schemes, similar to ARC, have been proposed and studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted. The robustness guarantees in (Gupta & Vaidya, 2020; Liu et al., 2021) only apply to strongly convex loss functions, under a specific redundancy condition (namely, $2f$-redundancy). We consider non-convex loss functions, as well as a generic heterogeneous setting of $(G,B)$-gradient dissimilarity. In contrast to the clipping scheme proposed in Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f/n$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients. Moreover, we prove a deterministic robustness property of ARC (see Theorem 3.2), whereas Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024) only provide probability guarantees assuming the non-adversarial inputs to be i.i.d. with the distribution satisfying certain special properties.
@@ 7 Conclusion & Discussion
Appendix A Computational Complexity of ARC.
The algorithm starts by computing the norms of all input vectors, which takes $\mathcal{O}(nd)$ time. Note that this step is also required for static clipping. The sorting operation of the norms takes $\mathcal{O}(n\log(n))$ time. Note that if $n$ is large compared to ${k=\left\lfloor{2(f/n)(n-f)}\right\rfloor}$, one can use more efficient algorithms for finding the ${(k+1)}$-th largest element, such as quick-select Hoare (1961) which has an average-case complexity of $\mathcal{O}(n)$. Finally, clipping the vectors requires $\mathcal{O}(nd)$ time. Therefore, the total complexity of ARC is $\mathcal{O}(nd+n\log(n))$. Overall, the complexity is comparable to static clipping, given that the sorting (or search) is crucially dimension independent.

### v5

@@ 4.2 Performance Gains of ARC Over Robust-DSGD
ARC increases the breakdown point in adverse settings. Figure 1(a) of Section 1 shows that for $f\in\{1,...,5\}$, Robust-DSGD with CWTM and GM completely fails to learn, consistently yielding worst-case maximal accuracies close to 15%. This suggests that $f=1$ constitutes the breakdown point for these aggregations in extreme heterogeneity. However, composing them with ARC increases the breakdown point of these aggregations to $f=3$. Indeed, for $f=1$ and 2, ARC enables Robust-DSGD to achieve accuracies greater than 85% in the worst case. However, when $f\geq 3$, the performance degrades, although CWTM with ARC is still able to reach a satisfactory accuracy close to 70%. Moreover, even when the heterogeneity is not large ($\alpha=0.5$), ARC still produces a significant improvement when the fraction of adversarial workers increases in the system. Indeed, in Figure 4(b) , the performances of ARC and no clipping are comparable for $f\leq 3$. However, the improvement induced by ARC is much more visible when $f\geq 5$. Particularly when $f=7$, ARC enables CWTM and GM to reach accuracies greater than 80% in the worst case, whereas the same aggregations yield accuracies below 40% without clipping, indicating the raise of the breakdown point due to ARC. Plots for $\alpha=0.1$ and 1 convey the same observations in Appendix F.1.
@@
Improved robustness on CIFAR-10. We also conduct experiments on CIFAR-10 with $n=17$ and $f=1$. Table 1 shows the worst-case maximal accuracies achieved by ARC and no clipping for CWTM and GM. For $\alpha=0.2$, ARC consistently outputs accuracies greater than 67% for both aggregations, while no clipping yields lower accuracies (with a larger variance). For instance, GM achieves 41.2% on average, i.e., 26% less than its counterpart with ARC. In the more heterogeneous setting $\alpha=0.075$, ARC enables all aggregations to reach worst-case maximal accuracies close to 60% (with a small variance). However, executing the same methods without clipping significantly deteriorates the performance of Robust-DSGD, with GM achieving 16% in worst-case accuracy. This can also be seen in Figure 23(a), where FOE completely degrades the learning when ARC is not used. We further extend this experiment to a larger distributed system with 33 honest workers, as detailed in Appendix F.3, where we observe the same trends as reported here.
@@ 5.1 Improvement of convergence guarantees
In this section, we characterize the improvement induced by ARC over the lower bound recalled in Lemma 2.3. Specifically, we consider Algorithm 1 with aggregation $\mathbf{F}\circ\mathbf{ARC}$, i.e., in each learning step $t$, $R_{t}\coloneqq\mathbf{F}\circ\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\dots,\,{{g}^{\left({n}\right)}_{t}}\right)$. We establish in Lemma 5.1 a key property of ARC, crucial to derive the improved stationarity error of Robust-DGD with ARC in Theorem 5.2.
Lemma 5.1 (Bounded Output).
Let $\mathbf{F}$ be an $(f,\kappa)$-robust aggregation method. For any vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ and set $S\subset[n]$ such that $|S|=n-f$, $\left\lVert{\mathbf{F}\circ\mathbf{ARC}(x_{1},\dots,x_{n})}\right\rVert\leq\max_{i\in S}\left\lVert{x_{i}}\right\rVert$.
@@
In the following, we show that when the local gradients of the honest workers at the initial model ${{\theta}_{1}}$ are sufficiently small, then Robust-DGD with ARC overcomes the lower bound recalled in Lemma 2.3. We denote $\mathsf{BP}\coloneqq\frac{1}{2+B^{2}}$, $\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\text{ and }\Psi(G,B,\rho)\coloneqq 640\left(1+\frac{1}{B^{2}}\right)^{2}\left(1+\frac{B^{2}\rho^{2}}{G^{2}}\right)$,
@@
Influence of model initialization and data heterogeneity. Note that the smaller the bound on the initial gradients (i.e., the better the model initialization), the greater the improvement induced by ARC. Specifically, a decrease in $\zeta$ leads to a reduction in $\rho$, which subsequently lowers $\Psi(G,B,\rho)$. As a result, for a fixed value of $\xi$ (i.e., for a fixed ratio of adversarial workers), the condition $\xi\leq\boldsymbol{\upsilon}/\Psi(G,B,\rho)$ is satisfied for a smaller $\boldsymbol{\upsilon}$, thereby yielding a lower stationarity error. This theoretical deduction is empirically validated in Section 5.2. A similar improvement occurs when increasing data heterogeneity, while keeping all other factors unchanged. Specifically, as $G$ increases, $\Psi(G,B,\rho)$ decreases, allowing for a smaller reduction factor $\boldsymbol{\upsilon}$ and thus a larger improvement in performance. This trend is also validated empirically in Figures 3 and 4(a).
Influence of $\boldsymbol{f/n}$. Additionally, for a fixed model initialization, an increase of $f/n$ towards the $\mathsf{BP}$ (i.e., as $\xi\to 0$) allows the reduction factor $\boldsymbol{\upsilon}$ to become smaller, leading to a greater improvement, as evidenced in Figures 1(a) and 4(b). Indeed, we show in Theorem C.4 (a more complete version of Theorem 5.2) that ARC effectively increases the $\mathsf{BP}$ of ($f,\kappa$)-robust methods from $\frac{1}{2+B^{2}}$ to $\frac{1}{2}$, provided the initial honest gradients are bounded in norm. Lastly, we would like to recall that when the workers’ gradients at model initialization are arbitrarily large, the convergence guarantee of Robust-DGD with ARC reduces to that of classic Robust-DGD (see Section 3).
Practical scope of Theorem 5.2.
Since $\xi=\frac{\mathsf{BP}-f/n}{\mathsf{BP}}$ and $\upsilon\geq\Psi(G,B,\rho)\xi$, it follows that the improvement factor $\upsilon$ is at least $\Psi(G,B,\rho)\frac{\mathsf{BP}-f/n}{\mathsf{BP}}$. For ARC to induce an improvement, we require $\upsilon<1$, which reduces to having $f>n\mathsf{BP}\left(1-\frac{1}{\Psi(G,B,\rho)}\right)$. Since Theorem 5.2 assumes $f<n\mathsf{BP}$, the improvement occurs when $f\in\left(n\mathsf{BP}\left(1-\frac{1}{\Psi(G,B,\rho)}\right),n\mathsf{BP}\right)$. The length of this interval, i.e., $n\frac{\mathsf{BP}}{\Psi(G,B,\rho)}$, depends on the size of the system $n$. In small systems where $n<\Psi(G,B,\rho)$, the length of the interval is smaller than 1 since $\mathsf{BP}<\nicefrac{{1}}{{2}}$. Therefore, at most one value of $f$, i.e., the largest integer smaller than $n\mathsf{BP}$, can lie in the improvement interval. Conversely, in large-scale systems where $n\in\Omega\left(\Psi(G,B,\rho)\log{\Psi(G,B,\rho)}\right)$, the improvement interval expands significantly and can be much larger than 1, thereby demonstrating ARC’s effectiveness in large-scale distributed learning.
@@ 5.2 Influence of model initialization on empirical robustness
Figure 5: Worst-case maximal accuracies achieved by Robust-DSGD with CWTM, on MNIST ($\alpha=0.1$), with 10 honest workers and 1 Byzantine worker. The x-axis represents worsening model initialization.
Figures 1(b) and 5 illustrate the effect of model initialization on the performance of Robust-DSGD with and without ARC, evaluated on MNIST with 10 honest workers. We investigate two regimes of heterogeneity: extreme heterogeneity and $\alpha=0.1$, and consider $f=1$ adversarial worker. The experiment proceeds as follows: we begin with the default model parameters initialized by PyTorch (i.e., the same ones used in Section 4), which represent a well-chosen set of initial parameters. These parameters are then scaled multiplicatively by a factor $\mu$ where larger values of $\mu$ correspond to progressively worse initialization, and vary $\mu\in\{1,...,5\}$. The results show that, under well-initialized conditions ($\mu=1$), ARC significantly enhances the performance of Robust-DSGD, achieving a substantial improvement in worst-case maximal accuracy, particularly under extreme heterogeneity. In this regime, ARC boosts accuracy by about 70% compared to plain Robust-DSGD (Figure 1(b)). This ($\mu=1$) corresponds to the initialization conditions of the empirical results presented in Section 4. As $\mu$ increases (i.e., the initialization worsens), the performance of ARC-enhanced Robust-DSGD gradually declines, with noticeable degradation starting from $\mu=4$. Nevertheless, even at this point, ARC still offers a performance advantage over Robust-DSGD without clipping. By $\mu=5$, the performance of Robust-DSGD with ARC converges to that of plain Robust-DSGD, both achieving an accuracy of around 10%. Another key observation from Figures 1(b) and 5 is that the behavior of Robust-DSGD with ARC closely mirrors that of Byzantine-free DSGD. Both exhibit similarly poor performance when $\mu=5$, and their accuracies are comparable when $\mu\leq 3$, especially when $\alpha=0.1$. This suggests that ARC is particularly effective at exploiting good model initialization, similar to the performance of DSGD in the absence of Byzantine workers. In contrast, plain Robust-DSGD struggles to fully leverage well-initialized models, as evidenced by its consistently lower accuracy (around 20%) in Figure 1(b). These findings highlight the important influence of model initialization on the robustness of aggregation methods, and empirically validate our theoretical findings in Section 5.1.
@@ 6 Related Work
Gradient clipping is a well-known technique for tackling exploding gradients in deep learning (Goodfellow et al., 2016). It has been extensively analyzed in the centralized setting (Zhang et al., 2020b; a; Koloskova et al., 2023), with applications to differential privacy (Pichapati et al., 2019). Moreover, we would like to note that gradient clipping has also been shown useful to obtain tighter convergence guarantees for stochastic gradient-descent methods in the case where the gradients have heavy-tailed noise (Gorbunov et al., 2020; 2022; Danilova, 2023). The considered proof techniques, however, do not apply directly to our adversarial distributed learning setting. A similar study on the impact of clipping also exists for distributed settings (Zhang et al., 2022; Khirirat et al., 2023).
@@
While adaptive clipping schemes, similar to ARC, have been studied in the context of robust aggregation (Gupta & Vaidya, 2020; Liu et al., 2021; Diakonikolas et al., 2020; Abdalla & Zhivotovskiy, 2024), critical differences should be noted. The robustness guarantees in Gupta & Vaidya (2020); Liu et al. (2021) only apply to strongly convex loss functions, under the specific $2f$-redundancy condition. We consider non-convex loss functions, as well as a generic heterogeneous setting of $(G,B)$-gradient dissimilarity. In contrast to the clipping scheme proposed in Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024), ARC only uses the $f$ parameter to tune the clipping threshold and does not rely on any additional a priori knowledge of the distribution of honest workers’ gradients. Moreover, we prove a deterministic robustness property of ARC (see Theorem 3.2), whereas Diakonikolas et al. (2020); Abdalla & Zhivotovskiy (2024) only provides probability guarantees assuming the non-adversarial inputs to be i.i.d. with the distribution satisfying certain special properties.
Lastly, prior work (Guerraoui et al., 2021; Zhu & Ling, 2022; Allouah et al., 2023c; Choffrut et al., 2024; Allouah et al., 2024) has considered the challenge of privacy preservation alongside robustness in distributed learning. In this context, the gradients are assumed to have bounded norms in order to control the sensitivity of the algorithm, as is common in differential privacy (DP) (Abadi et al., 2016). This is often enforced in practice through static gradient clipping. Our findings suggest that an adaptive approach, such as ARC, may provide a more reliable alternative. However, a rigorous analysis of ARC in the context of DP remains an important avenue for future research.
@@ 7 Conclusion & Discussion
Acknowledgment.
This work was supported in part by the FNS Project *Controlling the Spread of Epidemics: A Computing Perspective* (200021_200477) and the FNS Project *TruBrain* (20CH21_218778 / 1). We also extend our gratitude to the anonymous reviewers of the ICLR 2025 conference for their valuable insights and feedback.
Appendix A Additional Details on ARC
This section provides further insights into the design and computational complexity of ARC, explaining its adaptive clipping strategy and its theoretical foundations. We also analyze its efficiency compared to static clipping.
A.1 Design of ARC
Lemma B.1 in Appendix B shows that $\mathbf{F}\circ\mathbf{Clip}_{C}$ is $(f,\tilde{\kappa})$-robust, provided that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$. Note that this condition is impossible to guarantee when using a fixed clipping threshold that does not depend on the input vectors. In order to ensure that $\left\lvert{S\setminus S_{c}}\right\rvert\geq 1$ for all subsets $S$ of size $n-f$, the clipping threshold $C$ should be large enough such that less than $n-f$ input vectors are clipped. This motivates the design of an adaptive clipping strategy. Accordingly, we propose to choose a clipping threshold such that the total number of clipped vectors is of the form $\left\lfloor{\lambda(n-f)}\right\rfloor$, where $\lambda<1$. Note that it is natural to clip more vectors as the fraction of adversarial workers $\nicefrac{{f}}{{n}}$ increases, to control the impact of Byzantine behavior. Therefore, we set $\lambda\coloneqq\zeta\frac{f}{n}$ where $0\leq\zeta\leq 2$. Since $\frac{f}{n}<\frac{1}{2}$ and $\lambda<1$, the total number of clipped vectors $\left\lfloor{\zeta\frac{f}{n}(n-f)}\right\rfloor<n-f$ for all $\zeta\in[0,2]$. This constitutes the underlying idea behind our adaptive clipping strategy ARC, presented in Algorithm 2. Notably, our theoretical results hold for all $\zeta\in[0,2]$. However, empirical evaluations show that setting $\zeta=2$ consistently provides the best performance, leading us to adopt this value in Algorithm 2 and throughout the paper.
A.2 Computational Complexity of ARC
The computational complexity of ARC is comparable to static clipping, with the primary difference being an additional sorting step. The overall complexity can be broken down as follows:
  1. 1.
    Norm computation: Computing the norms of all input vectors requires $\mathcal{O}(nd)$ time. (This step is also required in static clipping.)
  2. 2.
    Sorting step: Sorting the norms incurs an additional $\mathcal{O}(n\log n)$ time. If $n$ is large relative to $k=\left\lfloor{2(f/n)(n-f)}\right\rfloor$, an efficient selection algorithm such as quick-select (Hoare, 1961) can be used instead, reducing the complexity to $\mathcal{O}(n)$ in the average case.
  3. 3.
    Gradient clipping: Clipping vectors requires $\mathcal{O}(nd)$ time.
Thus, the total time complexity of ARC is:
$\mathcal{O}(nd+n\log n).$
Crucially, this complexity remains dimension-independent in the sorting step, making it scalable while maintaining robustness.